OpenStudy (anonymous):

simplify (2+x^2)/(1+x^2)

OpenStudy (anonymous):

oh rearry

OpenStudy (anonymous):

i cant do this

OpenStudy (anonymous):

i see lol well we can use long division right?

OpenStudy (anonymous):

oh ok nvm then

OpenStudy (anonymous):

1 rem 1

jimthompson5910 (jim_thompson5910):

there's not much you can do here, you can use long division to get \[\large \frac{2+x^2}{1+x^2}=1+\frac{1}{1+x^2}\] but that's not much of a simplification

OpenStudy (anonymous):

yeah thats right lol i just got the same answer thanks:)